OpenStudy (anonymous):

Write sin(5theta) in terms of sin(theta) and cos(theta)

OpenStudy (anonymous):

sketch the graph?

OpenStudy (anonymous):

have to solve in terms of sin(theta) and cos(theta) and also using the double angle stuff....

OpenStudy (anonymous):

I don't understand the question, it's not enough condition to do something but graph

OpenStudy (anonymous):

Sorry, I was not specifying: Write sin(5theta) in terms of sin(theta) and cos(theta)

OpenStudy (anonymous):

it's a long long story. but I can give you some formula, you can do it by yourself: sin (5theta) = sin (4theta + theta) = sin (4theta) cos (theta) + cos (4theta) sin (theta) and then, sin(4theta) = sin (2theta + 2 theta), the same process to replace sin(4theta) in terms of sin(theta).

OpenStudy (anonymous):

cos (4theta)= cos(2theta +2theta)= cos^2(2theta)+ sin^2(2theta)

OpenStudy (jotopia34):

You can look up the Sum and Difference Formulas andv Product to Sum Formulas to understand where they come from too

OpenStudy (anonymous):

and then, cos^(2theta)= (2cos^29theta)-1)^2. oh my god. I lost

OpenStudy (anonymous):

ye, jotopia34 is right.

OpenStudy (anonymous):

everything has formula, what you need is being patient

OpenStudy (anonymous):

i usually get to cos 4(theta) then stuff gets really confusing each time i try, a different answer seems to pop up :/

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

The biggest problem is that I dont know what the right answer is so i dont know when to stop.

OpenStudy (anonymous):

stop at sin and cos(theta) only, not sin or cos (2or 4theta)

OpenStudy (anonymous):

you can have sin^2 (theta) or cos^2(theta) , just 1theta.

OpenStudy (anonymous):

oh ok thanks mate

OpenStudy (anonymous):

yw